OpenStudy (anonymous):

Which is a true statement? A. 6.52 < 6.5 B. –2.8 > –3.5 C. –11.1 > 10.7 D. –7.75 < –8.45

OpenStudy (anonymous):

My guess is B

OpenStudy (amorfide):

okay so A 6.52 <6.5 this is not true as 6.52 is larger than 6.5, and the sign means less than so this is wrong B –2.8 > –3.5 this is saying -2.8 is larger than -3.5 this is true as you are counting down from 0 C –11.1 > 10.7 this is saying -11.1 is larger than 10.7 this is false as 10.7 is above 0, -11.1 is below 0 D –7.75 < –8.45 this is saying -7.75 is less than -8.45 this is false as you are counting down from 0 and -8 comes after -7

OpenStudy (anonymous):

so its b

OpenStudy (amorfide):

yes

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Which is an example of the commutative property? A. –8 + (–13) = –13 + (–8) B. 11 + (–14 + 5) = (11 + (–14)) + 5 C. 16 + (–16) = 0 D. –(22 + 1) = –22 + (–1)

OpenStudy (anonymous):

hmmmm. I think A but I am 75% sure